OpenStudy (unklerhaukus):

Solve the Differential Equation \[2(1-x^2)\frac{\text dy}{\text dx}-(1+x)y=(1-x^2)^{1/2}\]

OpenStudy (unklerhaukus):

\[2(1-x^2)\frac{\text dy}{\text dx}-(1+x)y=(1-x^2)^{1/2}\]

OpenStudy (unklerhaukus):

is it a Bernoulli equation/?

OpenStudy (anonymous):

Get dy/dx by itself. It does look like a Bernoulli.

OpenStudy (unklerhaukus):

\[\frac{dy}{dx}-\frac{(1+x)}{2(1-x^2)}y=\frac{(1-x^2)^{1/2}}{2(1-x^2)}\] \[\frac{dy}{dx}-\frac{(1+x)}{2(1-x^2)}y=\frac{(1-x^2)^{-1/2}}{2}\]

OpenStudy (unklerhaukus):

A Bernoulli Equation has the form\[y'+p(x)y=q(x)y^n\]

OpenStudy (unklerhaukus):

so \[p(x)=-\frac{(1+x)}{2(1-x^2)}\]\[q(x)=\frac{(1-x^2)^{-1/2}}{2}\]\[n=0\]

OpenStudy (unklerhaukus):

is this right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

so i substitute \[v=y^{1-n}=y\]

OpenStudy (anonymous):

Find the integrating factor now

OpenStudy (anonymous):

\[I.F=e^{\int\limits_{}^{} P(x)}\]

OpenStudy (unklerhaukus):

integrating factor is \[R(x)=e^{\int p(x)\text dx}=e^{\int-\frac{(1+x)}{2(1-x^2)}\text dx}\] \[=e^{-\frac 12\int\frac{(1+x)}{(1-x^2)}\text dx}\]

OpenStudy (anonymous):

Integration is real easy one

OpenStudy (unklerhaukus):

should i use a substitution now easy? i can not see how it is easy

OpenStudy (unklerhaukus):

or should break the integral into two parts?

OpenStudy (anonymous):

\[\frac{1+x}{1-x^2}=\frac{1+x}{(1+x)(1-x)}\]

OpenStudy (unklerhaukus):

ah . clever

OpenStudy (unklerhaukus):

integrating factor is \[=e^{-\frac {1}{2}\ln(1-x)}=(1-x)^{-1/2}\]

OpenStudy (unklerhaukus):

right?

OpenStudy (anonymous):

see it doesn't get any better

OpenStudy (unklerhaukus):

alrighty then , i should be able to solve from here i think. (my computers is about to run outta batteries) if i can get this question done , ill be back tomorrow, but i think ill be right from here) Thankyou @NotSObright

OpenStudy (unklerhaukus):

\[2(1-x^2)\frac{\text dy}{\text dx}-(1+x)y=(1-x^2)^{1/2}\] \[\frac{\text dy}{\text dx}-\frac{(1+x)}{2(1-x^2)}y=\frac{(1-x^2)^{1/2}}{2(1-x^2)}\] \[\frac{\text dy}{\text dx}-\frac{(1+x)}{2(1-x^2)}y=\frac{1}{2(1-x^2)^{1/2}}\] \[\frac{\text dy}{\text dx}-\frac 12\frac{(1+x)}{(1+x)(1-x))}y=\frac{1}{2(1-x^2)^{1/2}}\] \[\frac{\text dy}{\text dx}-\frac 12\frac{1}{(1-x)}y=\frac{1}{2(1-x^2)^{1/2}}\] integrating factor \[R(x)=e^{-\frac 12\int\frac{1}{(1-x)}\text dx}=e^{\frac 12 \ln(1-x)\text dx}=(1-x)^{1/2}\] \[\frac{\text d }{\text dx}\left({y(1-x)^{1/2}}\right)=\frac 12\frac{(1-x)^{1/2}}{(1-x^2)^{1/2}}\] \[\frac{\text d }{\text dx}\left({y(1-x)^{1/2}}\right)=\frac 12\left(\frac{1-x}{(1-x)(1+x)}\right)^{1/2}\] \[y(1-x)^{1/2}=\frac 12\int\frac 1{\sqrt{(1+x)}}\text dx\] \[\text{let }1+x=w\] \[{\text dx}={\text dw}\] \[y\sqrt{1-x}=\frac 12\int w^{-1/2}\text dw\] \[y\sqrt{1-x}=\frac 12 \frac{w^{1/2}}{1/2}+c\] \[y\sqrt{1-x}=w^{1/2}+c\] \[y\sqrt{1-x}=\sqrt{1+x}+c\]